Find the area below f(x)= -x^2+ 4x + 3 and above g(x) = -x^3 + 7x^2 - 10x + 5 over the interval 1 ≤ x ≤ 2.
Answer by htmentor(1343)

The area below f(x) and above g(x) is the integral of the difference of the two
functions.
f(x) - g(x) = x^3 - 8x^2 + 14x - 2
The indefinite integral of f-g is:
h(x) = x^4/4 - 8x^3/3 + 14x^2/2 - 2x. To compute the area we evaluate this function at the endpoints of the interval:
h(2) - h(1) = 4.0833.